# On the Approach Towards Equilibrium Through Momentum-Dependent Relaxation: Insights from Evolution of the Moments in Kinetic Theory

### v2

@@ II.2.4 Computational Details
For $\Lambda=0$, the matrix is diagonal ($\mathbf{C}(\Lambda)\sim I$) and eigenvalues are positive and give rise to a pure decay. However, for non-zero $\Lambda$, $\mathbf{C}(\Lambda)$ is off-diagonal and is similar to the left shift operator,
$S\rho_{n,l}(\tau)\sim\rho_{n-\Lambda,l}(\tau)\leavevmode\nobreak\ .$ (33)
The spectrum of the operator $\mathbf{C}(\Lambda)$ is no longer positive, and a late-time decay is no longer guaranteed. In Fig.(1) the evolution of temperature for two different values of $\Lambda\in\{0,0.125\}$ are given. We see that the temperature evolution diverges from the expected late-time decay behaviour for $\Lambda=0.125$.
Figure 1: The evolution of temperature for $\Lambda=0$ and $\Lambda=0.125$. We see a growth in temperature when $\tau/\tau_{R}\sim 10$
II.2.5 Constraints on moment evolution
To circumvent the truncation problem, we can enforce physical constraints on moment evolution. It is reasonable to assume that the solution is continuous in the exponent $\Lambda$ and as $\Lambda$ tends to zero, the solution converges to the momentum-independent solution. As the momentum-independent solution converges to local equilibrium, we expect that $\Lambda$ values close to zero should inherit this behavior. We ensure these properties by identifying the growth ($\alpha$) and decay ($\gamma$)terms in evolution,
$\alpha$ $=-\frac{1}{\tau}\left[\mathcal{P}(n,l)\rho_{n,l-1}+\mathcal{R}(n,l)\rho_{n,l+1}\right]+C_{n,l,\Lambda}\,,$ (34)
$\gamma$ $=-\frac{1}{\tau}\left[\mathcal{Q}(n,l)\rho_{n,l}\right]\,.$ (35)
Figure 2: The effect of $\sigma$ on the evolution of $\rho_{2,5}$ for $\Lambda=0$
For equilibrium initial conditions, $\alpha$ dominates at the early time and eventually decreases after the moments reach their maximum values (combined effect of ${\cal{P,R}}$ and collision term). After that $\gamma$ drives the moments towards zero. We define a cutoff parameter $\sigma=\alpha/\gamma$ that can be varied at the cost of computing for finer time steps. When $\sigma$ is less than or equal to a predefined value, we set $\alpha=0$. This ensures that the moments decay to zero at late times. In Fig.(2) we plot the dependence of the moment $\rho_{2,5}/\rho_{2,0}$ for $\Lambda=0$ on the tuning parameter $\sigma$. From this baseline, we choose the tuning parameter to be around $\sigma\sim[10^{-4},10^{-5}]$.
III Results and discussions
@@
In Fig.(3(a)) we plot the scaled moment $\rho_{2,1}/\rho_{2,0}(=(P_{L}-P_{T})/\epsilon)$ which is related to the pressure anisotropy of the system as a function of scaled time $\bar{\tau}=\tau/\tau_{R}$. The moments have been plotted for various values of $\Lambda$ and two different sets of initial $\eta/s$. We see that for larger $\Lambda$, expansion generates larger anisotropies before decaying to zero. The reason is that the momentum-dependent relaxation time taken here scales as the positive exponents of the single-particle energy $E_{p}^{\Lambda}$. So, particles at higher energy (momentum) and with larger $\Lambda$ values are expected to have longer relaxation times and a slower decay rate before restoring the equilibrium. The separation between the peak anisotropy values are also observed to increase with larger momentum dependence. We further observe that as $\Lambda$ varies, the system with a smaller initial $\eta/s$ shows a larger relative increase in anisotropy and hence a greater sensitivity to $\Lambda$.
@@
In Fig.(3(b)) we plot the scaled moment $\rho_{2,5}/\rho_{2,0}$ as a function of $\bar{\tau}$. Larger $l$ moments give us the information about the higher angular variation of the out-of-equilibrium distribution function. Like in the previous $\rho_{2,1}$ case, the peak anisotropy value is higher for increasing $\Lambda$ values. It also shows a larger separation in the peak values when $\Lambda$ is increased. This indicates a sharper sensitivity of large $l$ moments on $\Lambda$. The sensitivity of $l$ moments on $\Lambda$ is particularly pronounced for lower $\eta/s$.
@@ IV CONCLUSIONS AND OUTLOOK
Existing studies with boost invariance and momentum-independent RTA show that in the moment hierarchy, the indices corresponding to the ranks of the irreducible momentum tensor (which, under Bjorken symmetry, align with even degrees of the Legendre polynomial), i.e, the $l$ index in $\rho_{n,l}$ become coupled. However, the indices related to the energy exponents $n$ do not mix and gives rise to decoupled equations of motion for each $n$-th moment. The situation becomes much more involved with MDRTA being introduced in the kinetic equation. The different $n$ moments now become coupled via the $\Lambda$ parameter (which controls the exponent in the momentum dependence in $\tau^{E}_{R}(E_{p})$) of MDRTA. The coupled differential equations now show atypical behavior of late time growth ($\tau/\tau_{R}\gg 1$) instead of decaying to zero. But by setting a tuning parameter, we approximated the behavior for a late time decay.
@@
This behavior is the result of interplay of competing mechanisms. Different angular anisotropies gets coupled through the free streaming dynamics with a $1/\tau$ dependence. While simultaneously, the collision kernel which depends on temperature and shear viscosity to entropy density ratio ($\eta/s$) introduces an admixture of $n$ moments that acts in opposition. The balance between these mechanisms ultimately governs the observed evolution of moments.
To the best of our knowledge, this work represents the first comprehensive investigation utilizing microscopic momentum transfers directly within the kinetic equation governing moment evolution. Our approach provides new understanding of the dynamics of non-equilibrium systems with momentum dependent relaxation. The analysis opens up the room for some theoretical concepts as well, such as the atypical behavior of late time growth could be traced from the current form of momentum dependence taken in $\tau^{E}_{R}(E_{p})$ which has been adopted from the existing literature [31, 33]. A more involved momentum dependence in the relaxation rate, especially depending on particle energy scales, could be a line of investigation for future endeavors.

### v5

@@ III.3.1 Moment Closure
where $A_{m,l}$ are expansion coefficients and the series is truncated at $N$ and $L$. We note that the Grad expansion in (41) is not in general convergent[28]. However, the series is asymptotic and can give accurate results for lower-order $p$ moments and when the system is near equilibrium. In our work we used $N=5$. The results were not found to be sensitive to the truncation order.
If we know the coefficients $A_{m,l}$, we can find any moment by integrating over the distribution function. Using the definition of the moments in (6) we get the set of relations between the integer moments $\rho_{n,l}$ and $A_{m,l}$ via the vector equations (Appendix (B.2)),
$\vec{\rho}_{l}$ $=\textbf{H}(T,0)\vec{\text{A}}_{l}\,$, (42)
where $\textbf{H}(T,\Lambda)$ is an $N\times N$ matrix, $\vec{\text{A}}_{l}$ and $\vec{\rho}_{l}$ are column vectors of moments $\rho_{n,l}$ and coefficients $A_{m,l}$ constructed by fixing $l$ and varying $n$ (Detailed definitions are given in Appendix(B.2)). This set of linear equations can be inverted to get $\text{A}_{m,l}$
$\vec{\text{A}}_{l}$ $=\textbf{H}^{-1}(T,0)\vec{\rho}_{l}\,.$ (43)
So knowing any $N\times L$ moments completely determines the coefficients $\text{A}_{m,l}$. Using this we approximate the fractional moments $\rho_{n-\Lambda}$ as
$\vec{\rho}_{n-\Lambda}$ $=T^{\Lambda}\textbf{H}(T,-\Lambda)\textbf{H}^{-1}(T,0)\vec{\rho}_{n}\,$, (44)
which expresses the fractional moments as a linear combination of the integer moments. The vector form of the moment equations now reads,
$\frac{d}{d\tau}\vec{\rho}_{l}=-\frac{1}{\tau}\mathbf{F}\vec{\rho}_{l}-\frac{1}{\tau^{0}_{R}}\textbf{M}(T,\Lambda)\vec{\rho}_{l}\,.$ (45)
where $\textbf{M}(T,\Lambda)=T^{\Lambda}\textbf{H}(T,-\Lambda)\textbf{H}(T,0)$. One can verify numerically that M has positive definite eigenvalues for an appropriate truncation in $N$. An analytical proof of the positive definiteness of $\textbf{M}(T,\Lambda)$ is given in (B.2).
IV Results and discussions
@@
In Fig.(3(a)) we plot the scaled moment $\rho_{2,1}/\rho_{2,0}(=(P_{L}-P_{T})/\epsilon)$ which is related to the pressure anisotropy of the system as a function of scaled time $\bar{\tau}=T^{-\Lambda}(\tau/\tau_{R}^{0})$. The moments have been plotted for various values of $\Lambda$ and two different sets of initial $\eta/s$. We see that for larger $\Lambda$, expansion generates larger anisotropies before decaying to zero. The reason is that the momentum-dependent relaxation time taken here scales as the positive exponents of the single-particle energy $E_{p}^{\Lambda}$. So, particles at higher energy (momentum) and with larger $\Lambda$ values are expected to have longer relaxation times and a slower decay rate before restoring the equilibrium. The separation between the peak anisotropy values are also observed to increase with larger momentum dependence. We further observe that as $\Lambda$ varies, the system with a smaller initial $\eta/s$ shows a larger relative increase in anisotropy and hence a greater sensitivity to $\Lambda$.
@@
In Fig.(3(b)) we plot the scaled moment $\rho_{2,5}/\rho_{2,0}$ as a function of $\bar{\tau}$. Larger $l$ moments give us the information about the higher angular variation of the out-of-equilibrium distribution function. Like in the previous $\rho_{2,1}$ case, the peak anisotropy value is higher for increasing $\Lambda$ values. It also shows a larger separation in the peak values when $\Lambda$ is increased. This indicates a sharper sensitivity of large $l$ moments on $\Lambda$. The sensitivity of $l$ moments on $\Lambda$ is particularly pronounced for lower $\eta/s$. We see that the decay time of the moment increases for increasing values of $\Lambda$. This implies that the system approaches equilibrium significantly slower when the relaxation time has a larger dependence (larger $\Lambda$) on momentum.
@@ V CONCLUSIONS AND OUTLOOK
Existing studies with boost invariance and momentum-independent RTA show that in the moment hierarchy, the indices corresponding to the ranks of the irreducible momentum tensor (which, under Bjorken symmetry, align with even degrees of the Legendre polynomial), i.e, the $l$ index in $\rho_{n,l}$ become coupled. However, the indices related to the energy exponents $n$ do not mix and gives rise to decoupled equations of motion for each $n$-th moment. The situation becomes much more involved with MDRTA being introduced in the kinetic equation. The different $n$ moments now become coupled via the $\Lambda$ parameter (which controls the exponent in the momentum dependence in $\tau_{R}(p)$) of MDRTA. The coupled differential equations now show atypical behavior of late time growth ($\tau/\tau_{R}\gg 1$) instead of decaying to zero. But by setting a tuning parameter, we approximated the behavior for a late time decay.
@@
(iii) Increased anisotropy at intermediate times for larger $\Lambda$ implies a later isotropisation and therefore slower approach to local equilibrium.
This behavior is the result of interplay of competing mechanisms. Different angular anisotropies gets coupled through the free streaming dynamics with a $1/\tau$ dependence. While simultaneously, the collision kernel which depends on temperature and shear viscosity to entropy density ratio ($\eta/s$) introduces an admixture of $n$ moments that act in opposition. The balance between these mechanisms ultimately governs the observed evolution of moments.
To the best of our knowledge, this work represents the first comprehensive investigation utilizing microscopic momentum transfers directly within the kinetic equation governing moment evolution. Our approach provides new understanding of the dynamics of non-equilibrium systems with momentum dependent relaxation. The analysis opens up the room for some theoretical concepts as well, such as the atypical behavior of late time growth could be traced from the current form of momentum dependence taken in $\tau_{R}(p)$ which has been adopted from the existing literature [31, 33]. A more involved momentum dependence in the relaxation rate, especially depending on particle energy scales, could be a line of investigation for future endeavors.